OpenStudy (anonymous):

Will give medal!! Solve -4(sqrt x+2)+3=-1

OpenStudy (ivettef365):

so you have the equation as follows: -4 \(\sqrt{x+2}\) +3 = -1

OpenStudy (anonymous):

yes

OpenStudy (ivettef365):

first subtract 3 on both sides and you have -4 \(\sqrt{x+2}\) +3 -3 = -1 -3 -4 \(\sqrt{x+2}\) = -4

OpenStudy (anonymous):

okay... then what? Sorry I'm kinda clueless when it comes to math

OpenStudy (ivettef365):

then divide by -4 on both sides and you have -4/-4\(\sqrt{x+2}\) = -4/-4 \(\sqrt{x+2}\) = 1

OpenStudy (ivettef365):

now square both sides and you have {\(\sqrt{x+2}\))^2 = (1)^2 that will be equal to x+ 2 = 1 now just solve for x

OpenStudy (anonymous):

okay so x=-1

OpenStudy (ivettef365):

correct